# UCINet0: A Machine Learning based Receiver for 5G NR PUCCH Format 0

### v1

@@ IV-C Confusion Matrices and Column Charts
The confusion matrices in Figures 9 and 10 indicate the overall multi-label classification performance of the model. Figures 9 and 10 show the multi-label confusion matrix for the best ($\Delta=0$) and worst ($\Delta=4$) cases of 2020dB SNR respectively, for both simulated data and hardware-captured data. Multi-label classification can be thought of as a combination of several independent binary classifiers. Diagonally dominant confusion matrices indicate that each of the binary classifiers is well-trained. For a given SNR, as $\Delta$ increases, the model leans more towards false positives because the model (whose 25th input is $\tilde{N}_{UE}\in[N_{UE},N_{UE}+\Delta]$) is trying to predict $\tilde{N}_{UE}$ number of UEs. The value of $\Delta$ is always positive, meaning that the number of expected UEs, as indicated by L2, is always greater than or equal to the actual number of UEs transmitting on a given resource allocation. Hence, the model almost always over-predicts.
Figure 9: Multi-label confusion matrix at SNR = 2020 dB and $\Delta$ = 0 for (a) simulated test dataset and (b) hardware captured test dataset. Here, $\alpha$ represents the UCI specific cyclic shift.
@@
Figure 10: Multi-label confusion matrix at SNR = 2020 dB and $\Delta$ = 44 for (a) simulated test dataset and (b) hardware captured test dataset.
However, it should be noted that the multi-label confusion matrix does not show the exact details of misclassifications. To gain a deeper understanding of the mistakes made by the model, consider the matrices in Figures 11 to 14. For each instance of the received PUCCH signal, there are two characteristics to correct classification. The first characteristic is correctly identifying the number of multiplexed UEs ($N_{UE}$) embedded in the signal. Even though this is not an explicit output of the NN model, we believe that the model is learning to estimate the number of multiplexed UEs. Our belief is based on the fact that when the number of possible multiplexed UEs is not given as metadata to NN input, the accuracy drops significantly. The second characteristic of correct classification is derived from the explicit output of the NN model, which is the correct prediction of the value of $\alpha$ for each of the multiplexed UEs.
Figures 11 and  12 show confusion matrices for the number of multiplexed UEs for both simulated data and hardware-captured data at an SNR of 0 dB and $\Delta$ values of 0 and $4$ respectively. Figures 13 and  14 confusion matrices for the number of multiplexed UEs for both simulated data and hardware-captured data at an SNR of $20$ dB and $\Delta$ values of 0 and $4$ respectively. We notice that all the confusion matrices for the prediction of the number of multiplexed UEs are largely diagonally dominant. Furthermore, for any SNR, when $\Delta=0$, the error in the predicted number of UEs fluctuates between +1 and -1 in the majority of instances, confirming the high accuracy values in Figure 7. As $\Delta$ increases, the error tends to grow on the positive side, meaning that the NN model predicts more UEs than actually present. As stated above, the overprediction is because $\Delta$ is always positive.
@@
Figure 11: Confusion matrix for the number of multiplexed UEs $N_{UE}$, at SNR = 0 dB and $\Delta$ = 0 for (a) simulated test dataset and (b) hardware captured test dataset.
@@
Figure 12: Confusion matrix for the number of multiplexed UEs $N_{UE}$, at SNR = 0 dB and $\Delta$ = 44 for (a) simulated test dataset and (b) hardware captured test dataset.
@@
Figure 13: Confusion matrix for the number of multiplexed UEs $N_{UE}$, at SNR = 2020 dB and $\Delta$ = 0 for (a) simulated test dataset and (b) hardware captured test dataset.
@@
Figure 14: Confusion matrix for the number of multiplexed UEs $N_{UE}$, at SNR = 2020 dB and $\Delta$ = 44 for (a) simulated test dataset and (b) hardware captured test dataset.
@@
Figure 15: Bar graph for the cyclic shift $\alpha$ at SNR = 0 dB, $\Delta$ = 44, and with $N_{UE}=\{0,1,2,\dots,12\}$ for (a) simulated test dataset and (b) hardware captured dataset.
Figure 16: Bar graph for the cyclic shift $\alpha$ at SNR = 2020 dB, $\Delta$ = 44, and with $N_{UE}=\{0,1,2,\dots,12\}$ for (a) simulated test dataset and (b) hardware captured dataset.
IV-D Implications of Model Performance on a 5G System
In general, both false detections and missed detections are detrimental to PUCCH Format 0 performance. However, from the context of sustaining a communication link between a gNB and a UE, a small number of false detections are more tolerable than missed detections. For example, when a gNB falsely detects a Scheduling Request, it allocates a small uplink bandwidth to the UE. Since the UE does not have any valid data to transmit, it would populate the uplink packet with padding information and indicate its lack of data in the Buffer Status Report (BSR) [21]. No further uplink grants will be allocated to the UE.
@@ IV-D Implications of Model Performance on a 5G System
Hence, occupying a small uplink bandwidth (due to false detection of an SR) is preferable to occupying a large downlink bandwidth (due to missed detection of ACKs) since recovery from a Radio Link Failure (RLF) involves heavy signaling for the UE to reattach.
@@
As shown in Figures 10,  12 and  14, at low SNRs or high $\Delta$ values, the proposed UCINet0 tends to overpredict the number of UEs actually present, rather than underpredict. This shows that the model, even when it makes a wrong prediction, is good at avoiding scenarios leading to RLFs.
IV-E Model Complexity Analysis
@@
Figure 17 shows the number of trainable parameters for different FCN architectures. The number of trainable parameters increases with the number of layers and neurons. An increase in the number of trainable parameters leads directly to an increase in the latency incurred due to fetching the weights from memory and neural network computations. Current 5G systems function with fundamental time units on the order of hundreds of microseconds. To cater to this requirement, the processing of all the physical layer channels allocated to each user must be completed in a duration of the order of tens of microseconds. For example, in our 5G Testbed [7], we use a subcarrier spacing of 3030kHz, which translates to a slot duration of $500\mu s$. Hence, the physical layer processing of all the users allocated in that slot needs to be completed within this time frame. If 10 PUCCH Format 0 instances are allocated in an uplink slot, each of them should be processed in no more than $50\mu s$. This time constraint also applies to AI/ML-based approaches such as neural networks. Hence, FPGAs are a promising choice for hardware implementation of such time-constrained neural networks.
@@
In order to deploy a version of UCINet0 on hardware devices such as FPGAs, the trade-offs between model performance (Figure 5) and model complexity (Figure 17) must be considered. These two parameters help us choose the best-suited model based on the available memory and compute resources on the hardware. The memory and bandwidth requirements for storing and retrieving the weights of Neural Networks like UCINet0 can be met by Xilinx’s latest Versal series FPGAs [22], which contain AI engines to perform matrix multiplications efficiently.
@@
Figure 17: Line graph showing the number of trainable parameters for different FCN architectures. These parameters increase with the number of layers and neurons.
@@ Conclusion
In this paper, we have designed a generalized Machine Learning based PUCCH Format 0 receiver that decodes all the combinations of UCI bits across several multiplexed users. The proposed UCINet0 model is a multi-label classifier that showed superior performance compared to conventional correlation-based approaches by a significant margin. The inclusion of field datasets in the testing of the model ensures robustness and the ability of the network to work in a wide range of scenarios. In addition to standard AI/ML metrics like accuracy, we have also provided insights into the interpretability of the NN model through confusion matrices and column charts of detection percentages of individual classes.
@@ Acknowledgment
The authors would like to thank the Department of Telecommunications (DOT), India for funding the 5G Testbed project and the Ministry of Electronics and Information Technology (MeitY) for funding this work through the project ”Next Generation Wireless Research and Standardization on 5G and Beyond”.

### v2

@@ VII-C CONFUSION MATRICES AND COLUMN CHARTS
The confusion matrices in Figures 15 and 16 present the overall performance of the multi-label classifier model. Figures 15 and 16 show the multi-label confusion matrix for the best ($\Delta=0$) and worst ($\Delta=4$) cases of 2020dB SNR, respectively, for simulated data. Multi-label classification can be thought of as a combination of several independent binary classifiers. Diagonally dominant confusion matrices indicate that each of the binary classifiers is well-trained. For a given SNR, as $\Delta$ increases, the model leans more towards false positives because the model (whose 25th input is $N_{UE}\in[\tilde{N}_{UE},\tilde{N}_{UE}+\Delta]$) is trying to predict $\tilde{N}_{UE}$ number of UEs. The value of $\Delta$ is always positive, meaning that the number of expected UEs, as indicated by L2, is always greater than or equal to the actual number of UEs transmitting on a given resource allocation. Hence, the model almost always over-predicts rather than missing a user’s detection.
@@
Figure 15: Multi-label confusion matrix at SNR = 2020dB and $\Delta$ = 0 for simulated test dataset
@@
Figure 16: Multi-label confusion matrix at SNR = 2020dB and $\Delta$ = 44 for simulated test dataset
It should be noted that the multi-label confusion matrix does not show the exact details of misclassifications. To gain a deeper understanding of the mistakes made by the model, consider the matrices in Figures 17 to 20. For each instance of the received PUCCH signal, there are two characteristics of correct classification. The first characteristic is correctly identifying the number of multiplexed UEs ($\tilde{N}_{UE}$) embedded in the signal. Even though this is not an explicit output of the NN model, we believe that the model is learning to estimate the number of multiplexed UEs. Our belief is based on the fact that when the number of possible multiplexed UEs is not given as metadata to the NN input, the accuracy drops significantly. The second characteristic of correct classification is derived from the explicit output of the NN model, which is the correct prediction of the value of $\alpha$ for each of the multiplexed UEs.
Figures 17 and  18 show confusion matrices for the number of multiplexed UEs for simulated data at an SNR of 0 dB and $\Delta$ values of 0 and $4$, respectively. Figures 19 and  20 are confusion matrices for the number of multiplexed UEs for simulated data at an SNR of $20$ dB and $\Delta$ values of 0 and $4$, respectively. We notice that all the confusion matrices for the prediction of the number of multiplexed UEs are largely diagonally dominant. Furthermore, for any SNR, when $\Delta=0$, the error in the predicted number of UEs fluctuates between +1 and -1 in the majority of instances, confirming the high accuracy values in Figure 10. As $\Delta$ increases, the error tends to grow on the positive side, meaning that the NN model predicts more UEs than are actually present. As stated above, the overprediction is because $\Delta$ is always positive.
@@
Figure 17: Confusion matrix for the number of multiplexed UEs $\tilde{N}_{UE}$, at SNR = 0 dB and $\Delta$ = 0 for simulated test dataset.
@@
Figure 18: Confusion matrix for the number of multiplexed UEs $\tilde{N}_{UE}$, at SNR = 0 dB and $\Delta$ = 44 for simulated test dataset.
@@
Figure 19: Confusion matrix for the number of multiplexed UEs $\tilde{N}_{UE}$, at SNR = 2020 dB and $\Delta$ = 0 for simulated test dataset.
Figure 20: Confusion matrix for the number of multiplexed UEs $\tilde{N}_{UE}$, at SNR = 2020 dB and $\Delta$ = 44 for simulated test dataset.
@@
Figure 21: Column chart for the cyclic shift $\alpha$ at SNR = 0 dB, $\Delta$ = 44, and with $\tilde{N}_{UE}=\{0,1,2,\dots,12\}$ for simulated test dataset.
Figure 22: Column chart for the cyclic shift $\alpha$ at SNR = 2020 dB, $\Delta$ = 44, and with $\tilde{N}_{UE}=\{0,1,2,\dots,12\}$ for simulated test dataset.
VII-D IMPLICATIONS OF MODEL PERFORMANCE ON A 5G SYSTEM
In general, both false detections and missed detections are detrimental to PUCCH Format 0 performance. However, from the context of sustaining a communication link between a gNB and a UE, a small number of false detections are more tolerable than missed detections. For example, when a gNB falsely detects an SR, it allocates a small uplink bandwidth to the UE. Since the UE does not have any valid data to transmit, it would populate the uplink packet with padding information and indicate its lack of data in the Buffer Status Report (BSR) [20]. No further uplink grants will be allocated to the UE.
@@ VII-D IMPLICATIONS OF MODEL PERFORMANCE ON A 5G SYSTEM
Hence, occupying a small uplink bandwidth (due to false detection of an SR) is more tolerable compared to occupying a large downlink bandwidth (due to missed detection of ACKs), since recovery from a Radio Link Failure (RLF) involves heavy signaling for the UE to reattach.
@@
As shown in Figures 16,  18, and  20, at low SNRs or high $\Delta$ values, the proposed UCINet0 tends to overpredict the number of UEs actually present, rather than underpredict. This shows that the model, even when it makes a wrong prediction, is good at avoiding scenarios leading to RLFs.
VII-E MODEL COMPLEXITY, MEMORY AND LATENCY ANALYSIS
@@
Figure 23 shows the number of trainable parameters for different FCN architectures. The number of trainable parameters increases with the number of layers and neurons. An increase in the number of trainable parameters leads directly to an increase in the latency due to (1) fetching the weights from memory and (2) corresponding neural network computations. Current 5G systems function with fundamental time units in the order of hundreds of microseconds. To cater to this requirement, processing all the physical layer channels allocated to each user must be completed in a duration of tens of microseconds. For example, for a subcarrier spacing of 3030kHz, the slot duration is of $500\mu s$. Hence, the physical layer processing of all the users allocated in that slot needs to be completed within this time frame. This time constraint also applies to AI/ML-based approaches such as neural networks. Hence, FPGAs are a promising choice for hardware implementation of such time-constrained neural networks.
In order to deploy a version of UCINet0 on hardware devices such as FPGAs, the trade-offs between model performance (Figure 4) and model complexity (Figure 23) must be considered. These two parameters help us choose the best-suited model based on the available memory and compute resources on the hardware. For the inference of UCINet0 for various test scenarios, we use an NVIDIA A100 Tensor Core GPU. Recall that our NN consists of an input layer with 24 neurons, two hidden layers with 257 and 256 neurons respectively, and an output layer with 12 neurons. This model consists of $75532$ trainable parameters ($(24\times 256+256)+(257\times 256+256)+(256\times 12+12)$). Similarly, the number of trainable parameters is calculated for other architectures and shown in Figure 23.
Assuming each parameter is represented by single precision floating point format (FP32), i.e., 4 bytes per parameter, resulting in a model memory of 302.128302.128KB. This is the memory required to store the model’s weights and biases both during training and inference. Along with this, activation memory is used to store the outputs of each layer during the forward pass of the network. With a batch size of 512, the total activation memory consumed by the model is $512\times(25+256+257+256+12)\times 4\text{bytes}=1.65$MB. To perform an inference on 1000 instances on the A100, the computed average runtime latency per instance is $0.27$ ms which is sufficient for a 5G network with 3030KHz which has a slot duration of $0.5$ ms.
@@
The memory and bandwidth requirements for storing and retrieving the weights of neural networks like UCINet0 can be met by Xilinx’s latest Versal series FPGAs [21], which contain AI engines to perform matrix multiplications efficiently.
@@
Figure 23: Line graph showing the number of trainable parameters for different FCN architectures. The number of FCN layers and the number of neurons of each architecture are presented on the x-axis. The number of trainable parameters corresponding to these architectures increases with the number of layers and neurons.
@@ CONCLUSION
In this paper, we have designed a generalized Machine Learning based receiver for PUCCH Format 0 that decodes all the combinations of UCI bits across several multiplexed users. The proposed UCINet0 model is a multi-label classifier that demonstrates superior performance under both slow and fast fading channel scenarios, compared to conventional correlation-based approaches by a good margin. The inclusion of hardware-captured datasets in the testing of the model adds robustness and shows the ability of the network to work in a wide range of scenarios. In addition to standard AI/ML metrics like accuracy, we have also provided insights into the interpretability of the NN model through confusion matrices and column charts of detection percentages of individual classes.
@@ ACKNOWLEDGMENT
The authors would like to thank the Department of Telecommunications (DOT), India for funding the 5G Testbed project and the Ministry of Electronics and Information Technology (MeitY) for funding this work through the project “Next Generation Wireless Research and Standardization on 5G and Beyond”.